Consider a population of 10,000 cranes that is decreasing exponentially by 300 cranes per year. (Type out your work and report units.)

What is the intrinsic rate of increase, r?

A crane expert estimates that the minimum viable population size is 1000 cranes. If nothing changes, how long will it take for this population to reach the minimum viable population size?
Answer #1

For an exponentially decreasing population, the value of intrinsic rate of decrease will be negative

dN/dt = rN

-300 = r × 10000

r = - 0.03

To find out the number of years it will take to reach minimum viable population,

ln(n2/n1) = r(t2-t1)

2.303log(1000/10000). = -(0.03)× t

-(2.303) = -(0.03)× t

Therefore, t. = 2.303/0.03

  = 76.7

  

So, it will take 77 years to reach the minimum viable population.